# Effects of Magnetic Helicity on 3D Equilibira and Self-Organized States in KTX Reversed Field Pinch

### v1

@@ 1 Introduction
The multi-region relaxed magnetohydrodynamics (MRxMHD) is a new fluid model that combines the ideal MHD and Taylor relaxation [18]. MRxMHD considers the helicity constraint of different subvolumes instead of the global constraint. Dennis et al. pointed out that the use of this model requires only a finite number of constraints to describe the self-organized plasma states in RFP [19]. However, this has not been fully investigated, and the impact of these constraints on the self-organized state of plasma has not been shown.
@@ 2 Multi-region relaxed magnetohydrodynamics
where the bracket [[*]]delimited-[]delimited-[][[*]][ [ * ] ] indicate the jump of the scalar across interface $\mathcal{I}$.
@@ 3 Self-organized equilibrium simulation of KTX
Figure 3: Poincaré cross section at $\zeta=0$ of the magnetic field with different $\mathcal{H}\right>$. No Poincaré plot in the outer subvolume because the toroidal field is very small in that region.
@@
The input data in these simulations are toroidal flux $\Psi_{i}$, stepped pressure $p_{i}$ and helicity $\mathcal{H}_{i}$, where $i=1,2$ represent the inner and outer subvolume, respectively. The difference between the two stepped-pressures is set to $p_{1}-p_{2}=560\ \mathrm{Pa}$ and the total toroidal flux is $\Psi=\Psi_{1}+\Psi_{2}=0.1298\ \mathrm{Wb}$ according to the discharge and design parameters of KTX. To keep the barrier in a relatively out position, the magnetic flux of the inner subvolume is $\Psi_{1}=0.86\Psi$. The SPEC outputs are the shape of the interface and the vector potential $\mathbf{A}$. The magnetic field can be obtained by $\mathbf{B}=\nabla\times\mathbf{A}$.
@@ 4 Magnetic field properties of different self-organized states
The magnetic shear of the new flux tube in the DAx state and the core in the SHAx state will vanish. Figure 7 shows the safety profiles of three self-organized states with different $\mathcal{H}\right>$. Both the DAx and SHAx states have lower safety factors than the axisymmetric state. The safety factor of the magnetic island in the DAx state and the core in the SHAx state is a constant of $1/6$. This shows that the helical magnetic field in the core of the SHAx state is a large magnetic island.
@@
There is an apparent positive linear correlation between $\mathcal{H}\right>$ and the plasma current. In different self-organized states, the current is enhanced with increasing magnetic helicity. Figure 8 shows this relationship. For KTX, when the average toroidal magnetic field $B_{t}\right>\approx 0.26\ \mathrm{T}$, we estimate that a plasma current of about $0.6\ \mathrm{MA}$ is needed to spontaneously transform into a three-dimensional state and when the current reaches $0.72\ \mathrm{MA}$, the plasma can self-organize to form the SHAx state. There is also a higher probability of obtaining QSH spectra with higher plasma current experimentally [1, 30, 31, 32].

### v2

@@ 1 Introduction
The multi-region relaxed magnetohydrodynamics (MRxMHD) is a new fluid model that combines the ideal MHD and Taylor relaxation [18]. MRxMHD considers the helicity constraint of different subvolumes instead of the global constraint. Dennis et al. pointed out that the use of this model requires only a finite number of constraints to describe the self-organized plasma states in RFP [19]. However, this has not been fully investigated, and the impact of these constraints on the self-organized state of plasma has not been shown. A multiple domains scheme (MDS) is proposed to investigate the transport problem in the presence of magnetic islands, which also divides the plasma into multiple regions [20]. The MDS was applied to RFP plasma, which has been implemented in a new numerical tool [21].
@@ 2 Multi-region relaxed magnetohydrodynamics
where the bracket [[]]delimited-[]delimited-[][[*]][ [ ∗ ] ] indicate the jump of the scalar across interface $\mathcal{I}$.
@@ 3 Self-organized equilibrium simulation of KTX
Figure 3: Poincaré plot at $\zeta=0$ in the inner subvolume with different $\mathcal{H}\right>$. The outmost red curve is the cross section of the interface.
@@
The input data in these simulations are toroidal flux $\Psi_{i}$, stepped pressure $p_{i}$ and helicity $\mathcal{H}_{i}$, where $i=1,2$ represent the inner and outer subvolume, respectively. The difference between the two stepped-pressures is set to $p_{1}-p_{2}=560\ \mathrm{Pa}$ and the total toroidal flux is $\Psi=\Psi_{1}+\Psi_{2}=0.1298\ \mathrm{Wb}$ according to the discharge and design parameters of KTX. The step of the pressure profile is determined by the magnetic flux of the subvolume. Based on previous results [4, 12], the transport barrier locates at a relatively outer position in the radial direction. In our simulaton, the flux of inner subvolume is $\Psi_{1}=0.86\Psi$, which keeps the position of the interface relatively out. The SPEC outputs are the shape of the interface and the vector potential $\mathbf{A}$. The magnetic field can be obtained by $\mathbf{B}=\nabla\times\mathbf{A}$.
@@ 4 Magnetic field properties of different self-organized states
The magnetic shear of the new flux tube in the DAx state and the core in the SHAx state will vanish. Figure 7 shows the inner subvolume safety factor profiles of three self-organized states with different $\mathcal{H}\right>$. Both the DAx and SHAx states have lower safety factors than the axisymmetric state. The safety factor of the magnetic island in the DAx state and the core in the SHAx state is a constant of $1/6$. This shows that the helical magnetic field in the core of the SHAx state is a large magnetic island.
@@
There is an apparent positive linear correlation between $\mathcal{H}\right>$ and the plasma current. In different self-organized states, the current is enhanced with increasing magnetic helicity. Figure 8 shows this relationship. The helicity flux density measurement system is implemented in KTX, and a similar linear relationship was obtained [28, 32]. For KTX, when the average toroidal magnetic field $B_{t}\right>\approx 0.26\ \mathrm{T}$, we estimate that a plasma current of about $0.6\ \mathrm{MA}$ is needed to spontaneously transform into a three-dimensional state and when the current reaches $0.72\ \mathrm{MA}$, the plasma can self-organize to form the SHAx state. There is also a higher probability of obtaining QSH spectra with higher plasma current experimentally [1, 33, 34, 35].